Three houses are available in a locality. Three persons apply for the houses. Each applies for one house without consulting others. The probability that all the three apply for the same house is ?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
For a particular house being selected
Probability = 1/3
Prob(all the persons apply for the same house) =(1/3)(1/3)(1/3)=1/9
Thanks and Regards
Shaik Aasif
askIITians faculty